# Equivalent class of Emergent Single Weyl Fermion in 3d Topological States: gapless superconductors and superfluids Vs chiral fermions

### v1

@@ VII.2 Charge for $\mathcal{T}$-symmetry invariant Hamiltonian
However, as in the first case, all the six linearly independent charge operators are dispersive in the momentum space $\mathcal{T}^{3}$ and can also contain the null operator. Charge operators in
$\{Q_{\alpha}({\bf k}),{\bf k}\in\mathcal{T}^{3};\alpha=1,...,6\}$ (96)
@@ VIII Conclusions
To summarize, we have put forward a general approach towards lattice chiral fermions starting with a gaped fermionic SPT which is time reversal invariant but breaks the charge $U(1)$ symmetry, i.e. a topological superconductor protected by the time reversal symmetry.
@@
III) The family of chiral lattice models with single Weyl fermions that have been known to us so far can be encoded in two dual copies of $3\otimes 3$-dimensional linear representations, or more precisely can be encoded in two dual $(1,1)$ representations of a Spin(4) group where one of the $SU(2)$ subgroups can also be identified as a subgroup of the emergent Lorentz group $SO(3,1)$. They form an equivalent class. In the infrared limit, they are isomorphic to either a tQCP with time reversal symmetry or its dual, a $T$-symmetry breaking superconducting nodal point phase.
IV) Furthermore, it is further utilized to pinpoint an intimate connection between a three spatial dimensional lattice chiral fermion model recently constructed in [53], with exact non-on-site non-compact symmetries, and real fermions that naturally appear in gapless superfluids or superconductors studied previously. We found that the proposal in the above reference is equivalent to applying a $T$-breaking action to a multiple-critical $T$-symmetry protected tQCP where the change of topologies, $\delta N_{w}=8$. This quadruples the more generic tQCP fundamental value of $\delta N_{w}=2$. This recent construction above is one example along Path c), i.e. a hydrid approach of Path a) and Path b).
@@
i) The conserved charge operators can span either a two-dimensional or a six-dimensional linear space at each $\bf p$}-point in the three torus $\mathcal{T}^{3}$. The conserved charge construction in this article as well as in previous chiral fermion discussions doesn’t appear to be unique. The role played by these large linear spaces and how they can be systematically and explicitly patched together in the space of $\mathcal{T}^{3}$ remain to be further studied.
@@ Appendix A Mathematical demonstration of the No-go theorem
Let $M=T^{3}\times\text{Herm}(n)$ and $H_{0}:T^{3}\rightarrow\text{Herm}(n)$.
We set $S:=\{(p,H(p))\mid p\in T^{3}\}$, which is clearly a 3-dimensional submanifold of $M$.
@@
We will see that $X=T^{3}\times\Sigma$ is a submanifold of $M$, and that although it is not closed (but $S$ is) it allows for a similar definition and properties of the intersection number as in III.1.
@@
We defined $X=T^{3}\times\Sigma$. Both $X$ and $S$ are orientable manifolds of complementary dimension, and as discussed in the main text, the intersection number $\text{int}(S,X)$ counts the sum of contributions from all Weyl cones in the Brillouin zone.
Consider a homotopy $F\colon[0,1]\times T^{3}\rightarrow M$ defined by:
@@
$\text{int}(S,X)=\text{int}(T^{3}\times\{A\},X)=0,$
@@ Appendix B Intersection number is invariant by homotopy
Assume that some point $x=(\lambda_{0},\mathbf{k}_{0})\in[0,1]\times T^{3}$ is in $\Sigma_{>2}$: then from dimensional considerations there is always a vector $\vec{v}$ orthogonal to $T_{x}(T^{3}\times\Sigma_{>2})$ and $T_{x}F([0,1]\times T^{3})$ to "push" this point out of $\Sigma_{>2}$ (by using a bump function $B$ that is non-zero only in the vicinity of 0 for example: $\tilde{F}(\lambda_{0}+\lambda,\mathbf{k}_{0}+\mathbf{k})=F(\lambda_{0}+\lambda,\mathbf{k}_{0}+\mathbf{k})+B(\lambda,\mathbf{k})\vec{v}$), by the orthogonality property and pythagoras theorem, this strictly adds distance between $S$ and $\Sigma_{>2}$.
@@ Appendix E Applying the Schrieffer-Wolff transformation
where $S_{k}$ is an hermitian operator chosen such that the transformed Hamiltonian $H^{\prime}(k)$ is block-diagonal to the desired order in perturbation theory.

### v2

@@ VII.2 Charge for $\mathcal{T}$-symmetry invariant Hamiltonian
However, as in the first case, all the six linearly independent charge operators are dispersive in the momentum space $\mathbb{T}^{3}$ and can also contain the null operator. Charge operators in
$\{Q_{\alpha}({\bf k}),{\bf k}\in\mathbb{T}^{3};\alpha=1,...,6\}$ (97)
@@ VIII Conclusions
To summarize, we have put forward a general approach towards lattice chiral fermions starting with a gapped fermionic SPT which is time reversal invariant but breaks the charge $U(1)$ symmetry, i.e. a topological superconductor protected by the time reversal symmetry.
@@
III) The family of chiral lattice models with single Weyl fermions that have been known to us so far can be encoded in two dual copies of $3\otimes 3$-dimensional linear representations, or more precisely can be encoded in two dual $(1,1)$ representations of a Spin(4) group where one of the $SU(2)$ subgroups can also be identified as a subgroup of the emergent Lorentz group $SO(3,1)$. They form an equivalence class. In the infrared limit, they are isomorphic to either a tQCP with time reversal symmetry or its dual, a $T$-symmetry breaking superconducting nodal point phase.
IV) Furthermore, it is further utilized to pinpoint an intimate connection between a three spatial dimensional lattice chiral fermion model recently constructed in [21], with exact non-on-site non-compact symmetries, and real fermions that naturally appear in gapless superfluids or superconductors studied previously. We found that the proposal in the above reference is equivalent to applying a $T$-breaking action to a multiple-critical $T$-symmetry protected tQCP where the change of topologies, $\delta N_{w}=8$. This quadruples the more generic tQCP fundamental value of $\delta N_{w}=2$. This recent construction above is one example along Path c), i.e. a hybrid approach of Path a) and Path b).
@@
i) The conserved charge operators can span either a two-dimensional or a six-dimensional linear space at each $\bf p$}-point in the 3-torus $\mathbb{T}^{3}$. The conserved charge construction in this article as well as in previous chiral fermion discussions doesn’t appear to be unique. The role played by these large linear spaces and how they can be systematically and explicitly patched together in the space of $\mathbb{T}^{3}$ remain to be further studied.
@@ Appendix A Mathematical demonstration of the No-go theorem
Let $M=\mathbb{T}^{3}\times\text{Herm}(n)$ and $H_{0}:\mathbb{T}^{3}\rightarrow\text{Herm}(n)$.
We set $S:=\{(p,H(p))\mid p\in\mathbb{T}^{3}\}$, which is clearly a 3-dimensional submanifold of $M$.
@@
We will see that $X=\mathbb{T}^{3}\times\Sigma$ is a submanifold of $M$, and that although it is not closed (but $S$ is) it allows for a similar definition and properties of the intersection number as in III.1.
@@
We defined $X=\mathbb{T}^{3}\times\Sigma$. Both $X$ and $S$ are orientable manifolds of complementary dimension, and as discussed in the main text, the intersection number $\text{int}(S,X)$ counts the sum of contributions from all Weyl cones in the Brillouin zone.
Consider a homotopy $F\colon[0,1]\times\mathbb{T}^{3}\rightarrow M$ defined by:
@@
$\text{int}(S,X)=\text{int}(\mathbb{T}^{3}\times\{A\},X)=0,$
@@ Appendix B Intersection number is invariant by homotopy
Assume that some point $x=(\lambda_{0},\mathbf{k}_{0})\in[0,1]\times\mathbb{T}^{3}$ is in $\Sigma_{>2}$: then from dimensional considerations there is always a vector $\vec{v}$ orthogonal to $T_{x}(T^{3}\times\Sigma_{>2})$ and $T_{x}F([0,1]\times\mathbb{T}^{3})$ to "push" this point out of $\Sigma_{>2}$ (by using a bump function $B$ that is non-zero only in the vicinity of 0 for example: $\tilde{F}(\lambda_{0}+\lambda,\mathbf{k}_{0}+\mathbf{k})=F(\lambda_{0}+\lambda,\mathbf{k}_{0}+\mathbf{k})+B(\lambda,\mathbf{k})\vec{v}$), by the orthogonality property and pythagoras theorem, this strictly adds distance between $S$ and $\Sigma_{>2}$.
@@ Appendix E Applying the Schrieffer-Wolff transformation
where $S_{k}$ is a Hermitian operator chosen such that the transformed Hamiltonian $H^{\prime}(k)$ is block-diagonal to the desired order in perturbation theory.